OpenStudy (anonymous):

CALCULUS HELP An Acura NSX going at 70 mph stops in 157 feet. Find the acceleration, assuming it is constant. I'm kind of confused, could I get a bit of help to push me towards the answer?

OpenStudy (amistre64):

assume it goes form 0 to 70 in 157 feet ... might be simpler

OpenStudy (amistre64):

and convertion from feet to miles might be useful

OpenStudy (welshfella):

I can solve this using the equations for uniform acceleration which are derived using calculus u = 70 v = 0 s = 157 a = ? v^2 = u^2 + 2as 0 = 70^2 + 314 a a = -(70^2) / 314

OpenStudy (anonymous):

I didn't want to use the kinetic equations, but I might have to.

OpenStudy (anonymous):

Sorry, kinematic

OpenStudy (amistre64):

let a = a v = at + k, v(0) = 0 soo v = at d = 1/2 at^2 + c, but d(0) = 0 sooo 157feet = 1/2 at^2 sqrt(2(157feet)/a) = t 70mph = a sqrt(2(157feet)/a) 70mph = sqrt(2a 157feet)

OpenStudy (welshfella):

i've forgotten the derivation i'm afraid now doubt Amistre will be able to help

OpenStudy (anonymous):

Ok, thanks. That makes sense.

OpenStudy (irishboy123):

\[\frac{dv}{dt} = -a\] does that mean anything to you?

OpenStudy (amistre64):

70mph^2 = 2a 157feet 70mph^2 --------- = a for whatever thats worth lol 2 *157feet

OpenStudy (anonymous):

I've already solved the problem...

OpenStudy (amistre64):

for a more general feel \[d_1=\frac12 at^2\] \[\sqrt{\frac{2d_1}{a}}=t\] \[v_f=a\sqrt{\frac{2d_1}{a}}\] \[v_f=\sqrt{2d_1~a}\] \[\frac{v_f^2}{2d}=a\]